check whether the function f (x,y)={4x^2y/4x^4+y^2 , (x,y)not=to (0,0) 0 (x,y)= to (0,0). } Is continuous at (0,0) ? Justify your answer.
1
Expert's answer
2020-04-06T13:00:36-0400
"f(x,y) = \begin{cases}
 \dfrac{4x^2y}{4x^4+y^2} &\text{if } (x,y)\not=(0,0) \\
 0 &\text{if } (x,y)=(0,0)
\end{cases}"

"\lim\limits_{(x,y)\to (0,0)}f(x,y)=[y=x^2]=\lim\limits_{(x,y)\to (0,0)}\dfrac{4x^2(x^2)}{4x^4+(x^2)^2}="

"={4\over 5}\not=0=f(0,0)"

Therefore the function "f(x,y)" is not continuous at "(0,0)."